OpenStudy (anonymous):

5x-4y+6=0 solve for y and find the slope. please show work, trying to figure out what i'm doing wrong.

OpenStudy (anonymous):

put it in the form y=mx+b

OpenStudy (anonymous):

m is the slope

OpenStudy (anonymous):

-4y=-5x-6. Getting y by itself. Y= -5/-4x-6/-4. Getting y by itself y=5/4x+3/2. Simplified. Slope is 5/4. Yintercept is 3/2

OpenStudy (anonymous):

Right. You're working for slope intercept form. y=mx+b

OpenStudy (anonymous):

DSS thank you very much

OpenStudy (anonymous):

So in other words we multiply the negative fractions times -1?

OpenStudy (anonymous):

A negative divided by a negative is a positive. -5/-4 = 5/4

OpenStudy (anonymous):

-6/-4 negative divided by negative is positive. 3/2. I think I missed typed part of my process

OpenStudy (anonymous):

no, I didn't. It justlooked confusing.

OpenStudy (anonymous):

Ok, so we treat every fraction as a division problem? I have another question how would you solve for y and x and the slope for 8(x+10y)=4

OpenStudy (texaschic101):

distribute through the parenthesis

OpenStudy (anonymous):

I did however the answers i'm inputting are not correct you get 8x+80y=4 you turn it into y=mx+b format and you get 80y=-8x+4 which would in turn give you y=10x+20 correct?

OpenStudy (texaschic101):

80y = -8x + 4 (now divide by 80) y = -8/80x + 4/80 y = - 1/10x + 1/20

OpenStudy (anonymous):

oh I see what I did wrong, I divided the denominator by the top number. thanks!

OpenStudy (texaschic101):

so the slope is -1/10 do you need to find the x and y intercept ?

OpenStudy (anonymous):

yeah I do

OpenStudy (anonymous):

question if I divided by 8 on the 8/80 I don't need to do it to the other fraction correct I can divide by 4?

OpenStudy (texaschic101):

to find the y intercept, sub in 0 for x 8x + 80y = 4 8(0) + 80y = 4 80y = 4 y = 4/80 y = 1/20 to find the x intercept, sub in 0 for y 8x + 80y = 4 8x + 80(0) = 4 8x = 4 x = 1/2 understand ?

OpenStudy (texaschic101):

yes, you can divide by 4, you do not need to divide by 8

OpenStudy (anonymous):

yes because we simplify after we divide. Thanks that makes a lot more sense then what the book explains. and to solve for x and y there is no need to put it into y=mx+b form correct?

OpenStudy (texaschic101):

not really, unless you are looking for just the y intercept, because in y = mx + b, the y intercept is in the b position

OpenStudy (texaschic101):

or if you are looking for the slope, it is in the m position

OpenStudy (anonymous):

oh ok, so with fractions do you a similar process, so for example a -5/3 x -7/4 y= 5

OpenStudy (texaschic101):

if you wanted to work this problem out without the fraction, multiply the equation by the common denominator, which is 12. -5/3x - 7/4y = 5 (multiply by 12) -20x - 21y = 60 now put it in y = mx + b -21y = 20x + 60 multiply by -1 to make x positive 21y = -20x - 60 divide by 21 y = -20/21x - 60/21 y = -20/21x - 20/7 or you can work with the fractions... -5/3x - 7/4y = 5 -7/4y = 5/3x + 5 y = (5/3x) / (-7/4) + (5)/(-7/4) y = 5/3x * -4/7 + 5 * - 4/7 y = - 20/21x - 20/7 either way, you will get the same answer

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

So whenever dividing a fraction you flip and multiply it?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

So the slope is -20/21 ?

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

and the y intercept is -20/7

OpenStudy (anonymous):

oh okay so I just tried doing the x intercept and lost myself again

OpenStudy (texaschic101):

sub in 0 for y

OpenStudy (anonymous):

so I get 20/21x=-20/7 correct?

OpenStudy (anonymous):

and now divide 20/21x by -20/7? which would turn into 21/20 * -20/7?

OpenStudy (texaschic101):

yes

OpenStudy (texaschic101):

yes 20/21x = -20/7 now divide by 20/21 x = (-20/7)/(20/21) x = -20/7 * 21/20

OpenStudy (texaschic101):

you flip what you are dividing by

OpenStudy (anonymous):

and you get 420/140 and we simplify it correct?

OpenStudy (anonymous):

which leaves me with 3?

OpenStudy (texaschic101):

don't forget about the sign -420/140 reduces to - 21/7 = -3

OpenStudy (anonymous):

Oh that's right forgot about the negative. Thank you very much! is there a way I can add you I'm starting my business major so if I have any further questions if you could possibly be of more help?

OpenStudy (texaschic101):

yes you can add me

OpenStudy (anonymous):

I don't know how.

OpenStudy (texaschic101):

click on my name and it will take you to my page and just click become a fan

OpenStudy (anonymous):

Got it thanks! have a great sunday.

OpenStudy (texaschic101):

you too :)